OpenStudy (anonymous):

The label on the car's antifreeze container claims to protect the car between −40°C and 125°C. To convert Celsius temperature to Fahrenheit temperature, the formula is C = 5 over 9 (F − 32). Write and solve the inequality to determine the Fahrenheit temperature range at which this antifreeze protects the car.

OpenStudy (anonymous):

easiest to first rewrite it as \[C = \frac{ 5 }{ 9 }(F - 32)\]\[\frac{ 9 }{ 5 }C = F - 32\]\[F = \frac{ 9 }{ 5 }C + 32\]

OpenStudy (anonymous):

plug in C = -40 to get the low F in the range then plug in C = 125 to get the high F in the range.

OpenStudy (anonymous):

-31/5

OpenStudy (anonymous):

thats what i got when i did 9/5 -40+32

OpenStudy (anonymous):

how did you get that? you want to plug in -40 then 125 to get a range: \[F = \frac{ 9 }{ 5 }(-40) + 32\]and\[F = \frac{ 9 }{ 5 }(125) + 32\]

OpenStudy (anonymous):

oh, you must mutliply :)

OpenStudy (anonymous):

oh ok give me a min

OpenStudy (anonymous):

-37 and 260

OpenStudy (anonymous):

i get -40 and 257

OpenStudy (anonymous):

pellet i did 35 not 32

OpenStudy (anonymous):

you are correct

OpenStudy (anonymous):

lol :)

OpenStudy (anonymous):

how do i got about choosing the answer choice

OpenStudy (anonymous):

the range in celcius was -40 to 125 the range in farenheit, by plugging in those values is -40 to 257 answer might look like this -40 F < T < 257 F

OpenStudy (anonymous):

−40 > 5 over 9 (F − 32) > 125; −40 > F > 257 −40 < 5 over 9 (F − 32) < 125; −40 < F < 257 −40 < 5 over 9 (F − 32); −40 < F 5 over 9 (F − 32) < 125; F < 257

OpenStudy (anonymous):

@divu.mkr

OpenStudy (anonymous):

second one